Can someone help me find the inverse of the following functions?? please???
f(x)=x+3
f(x)=x/(x+2)
f(x)=square root of x
f(x)=3^x....this one is 3 to the "x" power
please help me thank you!

Answer by stanbon(75887)

find the inverse of the following functions?? please???
f(x)=x+3
Write as y=x+3
Interchange x and y as follows:
x=y+3
Solve for "y", as follows:
y=x-3
That is the inverse.
--------------------------
f(x)=x/(x+2)
y=x/(x+2)
x=y/(y+2)
xy+2x = y
xy-y=-2x
y(x+2)=-2x
y=-2x/(x+2), x not equal to -2
That is the inverse.
--------------------------
f(x)=square root of x
y=sqrtx
x=sqrty
x^2=y
y=x^2, x>=0
That is the inverse.
-----------------------------
f(x)=3^x
y=3^x
x=3^y
y=log(base3)x
That is the inverse.
------------------
Cheers,
Stan H.